OpenStudy (anonymous):

Use the fundamental identities to simplify csc^2ß(α-cos^2ß)

OpenStudy (jdoe0001):

\(\bf \large csc^2(\beta)\cdot [\alpha-cos^2(\beta)]\quad ?\)

OpenStudy (anonymous):

what am I suppose to do?

OpenStudy (jdoe0001):

can you post a quick screenshot of the material?

OpenStudy (anonymous):

not really, it wont function properly.

OpenStudy (jdoe0001):

ok.... was thinking maybe something was missing

OpenStudy (anonymous):

no it just says use the fundamental identities to simplify

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ csc(\theta)=\cfrac{1}{sin(\theta)}}}\\ \quad \\ \quad \\ csc^2(\beta)\cdot [\alpha-cos^2(\beta)]\implies \left[\cfrac{1}{sin(\beta)}\right]^2\cdot [\alpha-cos^2(\beta)] \\ \quad \\ \cfrac{1^2}{sin^2(\beta)}\cdot [\alpha-cos^2(\beta)]\implies \cfrac{1}{sin^2(\beta)}\cdot [\alpha-cos^2(\beta)]\) distribute and simplify then

OpenStudy (anonymous):

Thank you so much

OpenStudy (jdoe0001):

yw